# Re(l)ality:

### v3

@@ 3 The View from Nowhere: Frame representations are perspectives on an equivalence class
Using the frame-bundle formalism introduced earlier, the model space is $M:=(\mathrm{Lor}(\mathcal{M})\cup\Phi)$, where $\Phi=\{\phi_{r}^{(I)},\phi_{b}^{(I)}\}$ represents the space of GPS scalars defining the red and the blue frames. This model space is structured as a principal fiber bundle with the structure group $\mathrm{Diff}(\mathcal{M})$ and base manifold $[M]$. Here, $[M]$ comprises equivalence classes of metrics and reference frame, expressed as $\{[m]\in M\}=\{[g_{ab}]\in[\mathrm{Lor}(\mathcal{M})],[\phi_{r/b}^{(I)}]\in[\Phi]\}$.
@@
Figure 2: The space of models $M=\mathrm{Lor}(\mathcal{M})\cup\Phi$ with its gauge group $\mathrm{Diff}(\mathcal{M})$. Each point corresponds to a triple $(g_{ab},\phi_{r}^{(I)},\phi_{b}^{(I)})$. A reference frame (either $\phi_{r}$ or $\phi_{b}$) selects a unique representative $g_{IJ}(\phi_{r})$ (or $g_{IJ}(\phi_{b})$) for each fiber $\mathbb{F}_{m}$. This is achieved through the projection map $f_{\phi_{r}}$ (or $f_{\phi_{b}}$). The map projects a model within a fiber onto the intersection of the fiber with a chosen reference frame, represented by the ‘level surface’ $F_{\phi_{r}}$ (or $F_{\phi_{b}}$).
As per Equation (2), $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ are gauge-invariant, relational observables that represent distinct physical scenarios. Nonetheless, they are frequently interpreted as mere perspectives on a shared equivalence class $[g_{ab}]$ of isometric metrics. This equivalence class is regarded as the sole structure with ‘ontological significance’, which means that it is considered the fundamental structure underlying physical reality. As shown in Table 1, $[g_{ab}]$ is frame-free, spatiotemporally implicit, and gauge-invariant. This interpretation constitutes what I call:
@@
Drawing from Kantian ontology (Kant, 1998), I differentiate between phenomenal reality, —measurable phenomena as they appear to observers— and noumenal reality, which represents the ultimate essence of reality. Within this framework, frame representations serve as phenomenal depictions of an implicit underlying reality, expressed through equivalence classes that cannot be explicitly characterised in spatiotemporal terms. In this terminology, the View from Nowhere presupposes the existence of a noumenal reality for which different spatiotemporal, phenomenological realities can be provided.
@@
Despite criticisms, the View from Nowhere maintains a substantial following. This is not surprising, since its conceptual foundations rest on what is known as the sophisticated approach to symmetries, or simply Sophistication about symmetries. This approach contrasts with Eliminativism, as it seeks to preserve symmetry-related structures rather than eliminate them entirely. For a more extensive analysis of the various positions on symmetries, see, e.g., Dewar (2019) and Gomes (2023b).
Eliminativism argues that a theory must be formulated solely in terms of symmetry-invariant objects such as equivalence classes. This eliminates individual symmetry-related models of the theory, thereby eradicating representational redundancy.
@@
In contrast, Sophistication rejects Eliminativism while embracing a structuralist stance. It posits that symmetries uncover an underlying invariant structure, which holds genuine ontological significance. Unlike Eliminativism, Sophistication allows symmetry-related models to coexist without removing them from the formalism. Jacobs (2021), drawing on the Erlangen Programme (Klein, 1893), characterises Sophistication as a symmetry-first approach, also known as the external approach (Dewar, 2019, p.502).
Moreover, Dewar argues that Sophistication, unlike Eliminativism, does not require an intrinsic characterisation of $[M]$ that makes no reference to the elements of $M$ for an ontological commitment to equivalence classes alone. However, Martens and Read (2020) have criticised this method of committing to invariant structures — defined as the interpretational approach by Møller-Nielsen (2017) — for failing to ensure a perspicuous understanding of the shared ontology underlying various models.
The interpretational approach asserts that two symmetry-related models can only be considered physically equivalent if a perspicuous characterisation of their shared ontology is provided.
In contrast, the motivational approach does not require perspicuity as a strict ontological condition but seeks justification for treating symmetry-related models as equivalent.
The ability to provide a perspicuous characterisation of the ontology of the theory aligns more naturally with the interpretational approach, whereas the motivational approach only requires to justify why symmetry-related models can be treated as equivalent.
@@
In response to this critique, Gomes (2024a) defends the use of reference frames (which he terms representational conventions), formalised through a projection operator $f_{\sigma}:M\to M$ on the fiber bundle $M$. This procedure provides a perspicuous, yet choice-dependent, characterisation of $[M]$. The adoption of $f_{\sigma}$ instead of the section map $\sigma:[M]\to M$ aligns with the Sophistication approach’s claim that intrinsic parameterisation of elements $[m]\in[M]$ is unnecessary. Furthermore, Gomes’ approach integrates the interpretational approach with the motivational approach, shielding Sophistication from accusations of being “cheap” (Martens and Read (2020); Belot (2017)) and providing a perspicuous understanding of the ontology. That is, Gomes succeeds in demonstrating that answering why symmetry-related models represent the same physical reality, as required by the motivational approach, also gives us an insightful characterisation of ontology as necessarily required by the interpretational approach.
@@
I argue that Gomes’ approach, while insightful, is ultimately insufficient. Even if one characterises each equivalence class by selecting a ’representative’ (a relational gauge-invariant observable), this does not resolve the underlying ontological commitment to the equivalence classes themselves. Designating a representative only provides a single perspective on the broader structure, leaving the full ontology of the theory incompletely (not perspicuously) addressed.
@@
I have already expressed many of my qualms about the View from Nowhere, and in the next section I will offer an alternative on how to interpret the formalism of relational observables. Before doing so, however, I want to dwell on a further problem related to the View from Nowhere: namely the loss of physical information. [^19] This loss is exemplified by what Tong says about the choice of a gauge: The [gauge] redundancy allows us to make manifest the properties of quantum field theories, such as unitarity, locality, and Lorentz invariance, that we feel are vital for any fundamental theory of physics but which teeter on the verge of incompatibility. If we try to remove the redundancy by fixing some specific gauge, some of these properties will be brought into focus, while others will retreat into murk. By retaining the redundancy, we can flit between descriptions as is our want, keeping whichever property we most cherish in clear sight. (Tong, 2018, p.1) It can be said that, within this view, different choices of reference frames can be conceptualised as ‘windows of knowledge that provide partial views of a shared invariant structure’, available to an observer (see Adlam (2024) on how to schematise a conscious observer in a diff-invariant theory such as GR). Accordingly, an observer can adjust her perspective to explore specific aspects of this invariant structure, effectively selecting one reference frame over another based on her investigative focus.
@@
This idea also resonates with Einstein (1917)’s assertion that “ […]a definite choice of the system of reference […] is contrary to the spirit of the relativity principle.” Similarly, (Adlam, 2024, p.9) argues that “diffeomorphism invariance could finally be broken by the observer herself.” Geng (2024) also echoes this viewpoint.
@@ 4 The View from Everywhere: Frame representations are all that exist
In Bamonti and Gomes (2024b), a map m called external diffeomorphism is introduced to relate the two frames $\{\phi_{r}^{(I)}\}$ and $\{\phi_{b}^{(I)}\}$, functioning analogously to a coordinate transformation. Unlike an ‘ordinary diffeomorphism’, an external diffeomorphism acts directly on the already constructed local, gauge-invariant observables, changing frames and getting us to a different and new observable.
@@
It is important to note that while there can indeed be frame-independent observables, these quantities, however, would inherently be non-local in nature (see Table 1).
@@
Also, the existence of the external diffeomorphism m enables advocates of the ”View from Everywhere” to define an equivalence class $[g_{IJ}]:=\{g_{IJ}(\phi_{r}),g_{IJ}(\phi_{b}),\dots\}$, which differs fundamentally from the equivalence class $[g_{ab}]:=\{g_{ab},(d^{*}g)_{ab},\dots\}$ used in the View from Nowhere. Unlike the latter, the equivalence class $[g_{IJ}]$ within the View from Everywhere supervenes on the ensemble of all the frame-dependent quantities. Each of these quantities form the a bona-fide physical reality, rather than existing as perspectives of an abstract, frame-free structure. Coherently, $[g_{IJ}]$ is a frame-independent object, as it is expressed across all possible reference frames. So differently from $[g_{ab}]$, it is not a frame-free object. In formal terms: the View from Everywhere retains a one-to-one correspondence between the equivalence class $[g_{IJ}]$ and the orbit $\mathcal{O}_{g}:=\{g_{IJ}(\phi_{r}),g_{IJ}(\phi_{b}),\cdots\}$. However, unlike the View from Nowhere, $[g_{IJ}]$ in the View from Everywhere does not exist independently of the frame-specific representations within the space of models. Essentially, $[g_{IJ}]$ and $\mathcal{O}_{g}$ coincide. See Figure 3.
@@
Figure 3: The set $[M]$ of equivalence classes is part of the space of models itself. Moreover, it is the set of relational observables that can be understood as a fiber, generated by the group $Diff(\mathbb{R}^{4})$, whose group elements are external diffeomorphisms m.
@@
If for the View from Nowhere the equivalence class $[g_{ab}]$ represented a shared ontology, for the View from Everywhere, the equivalence class $[g_{IJ}]$ only represents the presence of a shared vocabulary. This distinction underscores the divergence between the two views: the View from Everywhere rejects the existence of a frame-free, ontologically independent structure, focusing instead on the relational and frame-specific nature of observables. Consequently, the term ’perspective’ becomes problematic in this context, as it implies an underlying reality that can be described from multiple viewpoints—an assumption the View from Everywhere does not adopt.
@@
What conception of reality emerges from this framework? Within the View from Everywhere, each of the frame-dependent physical situations constitutes a fully-fledged reality. The collection of such realities constitutes what I call Relality. The concept of Relality is intriguing and deserves to be analysed further. Importabtly, this concept helps me to emphasise that reality is fundamentally local and relational.
Drawing on the Kantian perspective used to interpret the View from Nowhere, I argue that the View from Everywhere fundamentally rejects the notion of a shared, implicit noumenal reality underlying its phenomenological manifestations. According to the View from Everywhere, there is no ’hidden’ structure beyond these frame-specific representations; each gauge-invariant quantity constitutes the entirety of what fundamentally exists. Accordingly, the two observables $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ do not describe two distinct phenomenal perspectives of a shared, ontologically independent noumenal reality represented by $[g_{ab}]$. Instead, each constitutes a reality and there is nothing else. This is because the concept of Relality, formalised by $[g_{IJ}]$, does not represent a shared ontology. Its utility lies only in ensuring the inter-translatability of each reality. This supports the abandonment of the concept of ‘perspective’.
@@
In this section, I argue that the interpretive framework of the View from Everywhere challenges the core assumptions of the Sophistication approach. Specifically, I will demonstrate that the View from Everywhere offers greater ontological parsimony—minimising unnecessary metaphysical commitments—and provides a clearer, more perspicuous characterisation of the ontology underlying relational observables. Consequently, I contend that Sophistication is neither the ultimate nor the most effective framework for interpreting relational observables.
@@
Therefore, supporting the View from Nowhere does not imply the removal of redundant gauge degrees of freedom. Rather, it only involves attributing ontological importance to the gauge-invariant observables.
@@
First, the breaking of the diffeomorphic freedom of theory is precluded by the existence of the external diffeomorphism m. This map relates different reference frame choices232323m serves as the analogue of the transition map between sections of a principal bundle, given by $\mathrm{\bf{t}}_{\sigma\sigma^{\prime}}(\varphi)=f_{\sigma^{\prime}}(\varphi)^{-1}f_{\sigma}(\varphi)$ (see Gomes (2024a))., ensuring GR to remain m-covariant (see also Bamonti and Gomes (2024a)).
Second, the notion of Relality implies that all the available information resides within each frame. Consequently, there is no ‘total’ information that could be lost when a specific reference frame is selected. This may also alleviate the concerns raised in Wallace (2024) that the choice of certain reference frames (some gauge-fixings) does not preclude the physical features of the system from varying under a gauge transformation, leading to a loss of genuine information and casting doubt on whether they can really be classified as gauge-fixing at all (see ‘The Incompleteness Reason’, ivi, p.13). [^24] To be precise, Wallace studies the so-called ‘unitary gauge’ in electrodynamics. I argue that this choice of gauge corresponds to the use of uncoupled reference frames. In other words, the physical features that change are features of the physical system considered as isolated from the environment (Wallace, 2022a, b). This isolation procedure coincides with the dynamic uncoupling of the reference frame. The “loss of genuine information” referred to by Wallace, therefore, is not really genuine because it is a consequence of this approximate procedure. Within the View from Nowhere, on the other hand, there is the loss of bona-fide genuine information as a result of the use of coupled reference frames, which identify distinct gauge-invariant observables. The View from Nowhere alleviates this, more serious, problem.

### v4

@@ 3 The View from Nowhere: Frame representations are perspectives on an equivalence class
Using the frame-bundle formalism introduced earlier, the model space is $M:=(\mathrm{Lor}(\mathcal{M})\cup\Phi)$, where $\Phi=\{\phi_{r}^{(I)},\phi_{b}^{(I)}\}$ represents the space of GPS scalars defining the red and the blue frames. This model space is structured as a principal fibre bundle with the structure group $\mathrm{Diff}(\mathcal{M})$ and base manifold $[M]$. Here, $[M]$ comprises equivalence classes of metrics and reference frame, expressed as $\{[m]\in M\}=\{[g_{ab}]\in[\mathrm{Lor}(\mathcal{M})],[\phi_{r/b}^{(I)}]\in[\Phi]\}$.
@@
Figure 2: The space of models $M=\mathrm{Lor}(\mathcal{M})\cup\Phi$ with its gauge group $\mathrm{Diff}(\mathcal{M})$. Each point corresponds to a triple $(g_{ab},\phi_{r}^{(I)},\phi_{b}^{(I)})$. A reference frame (either $\phi_{r}$ or $\phi_{b}$) selects a unique representative $g_{IJ}(\phi_{r})$ (or $g_{IJ}(\phi_{b})$) for each fibre $\mathbb{F}_{m}$. This is achieved through the projection map $f_{\phi_{r}}$ (or $f_{\phi_{b}}$). The map projects a model within a fibre onto the intersection of the fibre with a chosen reference frame, represented by the ‘level surface’ $F_{\phi_{r}}$ (or $F_{\phi_{b}}$).
As per Equation (2), $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ are gauge-invariant, relational observables that represent distinct physical scenarios. Nonetheless, they are frequently interpreted as mere perspectives on a shared equivalence class $[g_{ab}]$ of isometric metrics. This equivalence class is regarded as the sole structure with ‘ontological significance’, which means that it is considered the fundamental structure underlying physical reality. [^22] To be blunt, this ontological picture, termed ‘Cartesian-Hegelian’ by Adlam (2024b), is not the same as that advocated by the moderate perspectivalism, which in contrast denies the claim that only the perspective-neutral picture, here represented by $[g_{ab}]$ is ontologically fundamental. According to moderate perspectivalism, “both perspectival and perspective-neutral facts exist and both are equally fundamental” [my emphasis] (ivi, p.9). However, my criticism of perspectivalism and the ontological fundamentality of $[g_{ab}]$ also extends to the moderate perspectivalism advocated by Adlam. In particular, I criticise the presence of the term ‘both’ on Adlam’s quote, in favour of the View From Everywhere, whereby equivalence classes $[g_{ab}]$ have no ontological status at all. Perspective-neutral facts exist in the same way that the number 5 exists. Note that the fact that moderate perspectivalism also recognises ontological fundamentality to the perspective-neutral structure supports the claim made in the introduction that moderate perspectivalism and the View From Nowhere are not at the antipodes. Both views support the ontological fundamentalism of equivalence classes. Remember the mantra: perspectives are always perspectives on something. As shown in Table 1, $[g_{ab}]$ is frame-free, spatiotemporally implicit, and gauge-invariant. This interpretation constitutes what I call:
@@
Drawing from Kantian ontology (Kant, 1998), I differentiate between phenomenal reality, —measurable phenomena as they appear to observers— and noumenal reality, which represents the ultimate essence of reality. Within this framework, frame representations serve as phenomenal depictions of an implicit underlying reality, expressed through equivalence classes that cannot be explicitly characterised in spatiotemporal terms. In this terminology, the View from Nowhere presupposes the existence of a noumenal reality for which different spatiotemporal, phenomenological realities can be provided. [^24] This noumenal reality resonates strongly with Adlam (2024b)’s claim that perspective-neutral reality is not empirically accessible (so, not phenomenological in Kantian’s terms), but it is there. Explicitly, she states: “Although these [perspective-neutral] facts do not directly describe possible experiences […] they are physically real” (ivi, p.7).
@@
Despite various criticisms, the View from Nowhere continues to enjoy a substantial following. Its conceptual basis is the so-called Sophistication approach to symmetries, which sharply contrasts with Eliminativism (or reductionism).
According to Eliminativism, a theory should be reformulated exclusively in terms of its symmetry-invariant quantities — typically equivalence classes — and thus eliminating all redundant, symmetry-related models.
@@
In contrast, Sophistication adopts a structuralist stance by retaining the full set of symmetry-related models and treating them as isomorphic representations of a single, underlying invariant structure, which holds genuine ontological significance. Indeed, as Jacobs (2021) (drawing on Klein’s Erlangen Programme (Klein, 1893)) explains, Sophistication is best characterised as a symmetry-first (or external) approach. For further discussion of these positions, see, e.g., Dewar (2019) and Gomes (2023b).
A central debate in this context concerns the justification for treating symmetry-related models as physically equivalent. Two interrelated strategies have been proposed (Møller-Nielsen, 2017).
The interpretational approach insists that physical equivalence can only be granted once a clear and explicit (i.e., perspicuous) account of the invariant ontology is provided.
In contrast, the motivational approach is less demanding ontologically and adopts a more pragmatic stance. Rather than requiring an exhaustive intrinsic description of the shared structure, it demands only a compelling justification for treating the models as equivalent.
In this sense, Eliminativism aligns with the interpretational approach by seeking to eliminate all non-invariant features, whereas Sophistication typically resonates with the motivational approach by justifying equivalence without necessitating full intrinsic characterisation of the invariant structure.
However, proponents of Sophistication, notably Dewar, contend that Sophistication does not require an intrinsic characterisation of the space $[M]$ (the collection of equivalence classes) in order to secure ontological commitment to its elements. He maintains that the symmetries themselves—by revealing the invariant structure—suffice for ontological commitment in line with the interpretational approach. However, this position has not gone unchallenged. Critics such as Martens and Read (2020) argue that the sophistication strategy, when understood in this light, is ‘cheap’ because it fails to provide a perspicuous account of the shared invariant ontology underlying the various models.
@@
In response to this critique, Gomes (2024a) defends Sophistication through the use of reference frames—termed representational conventions—which are formalised by a projection operator $f_{\sigma}:M\to M$ on the fibre bundle $M$. This procedure provides a perspicuous, yet choice-dependent, characterisation of $[M]$. The adoption of $f_{\sigma}$ in lieu of the section map $\sigma:[M]\to M$ supports the Sophistication claim that an intrinsic parametrisation of elements $[m]\in[M]$ is unnecessary. Furthermore, Gomes’ approach seeks to integrate aspects of both the interpretational and motivational approaches, shielding Sophistication from accusations of being ‘cheap’ and providing a perspicuous understanding of the ontology. That is, Gomes succeeds in demonstrating that answering why symmetry-related models represent the same physical reality, as required by the motivational approach, also gives us an insightful characterisation of ontology, as necessarily required by the interpretational approach.
@@
Nonetheless, I argue that Gomes’ approach, while insightful, is ultimately insufficient. Even if one characterises each equivalence class by selecting a ’representative’ (a relational gauge-invariant observable $f_{\sigma}^{*}(\bullet)$), this does not justify the underlying ontological commitment to the equivalence classes themselves. In effect, designating a representative only offers one perspective on the broader structure, leaving the full ontology of the theory inadequately (or insufficiently perspicuously) addressed.
@@
I have already expressed many of my qualms about the View from Nowhere, and in the next section I will offer an alternative on how to interpret the formalism of relational observables. Before doing so, however, I want to dwell on a further problem related to the View from Nowhere: namely the loss of physical information. [^26] This loss is exemplified by what Tong says about the choice of a gauge: The [gauge] redundancy allows us to make manifest the properties of quantum field theories, such as unitarity, locality, and Lorentz invariance, that we feel are vital for any fundamental theory of physics but which teeter on the verge of incompatibility. If we try to remove the redundancy by fixing some specific gauge, some of these properties will be brought into focus, while others will retreat into murk. By retaining the redundancy, we can flit between descriptions as is our want, keeping whichever property we most cherish in clear sight. (Tong, 2018, p.1) It can be said that, within this view, different choices of reference frames can be conceptualised as ‘windows of knowledge that provide partial views of a shared invariant structure’, available to an observer (see Adlam (2024a) on how to schematise a conscious observer in a diff-invariant theory such as GR). Accordingly, an observer can adjust her perspective to explore specific aspects of this invariant structure, effectively selecting one reference frame over another based on her investigative focus.
@@
This idea also resonates with Einstein (1917)’s assertion that “ […]a definite choice of the system of reference […] is contrary to the spirit of the relativity principle.” Similarly, (Adlam, 2024a, p.9) argues that “diffeomorphism invariance could finally be broken by the observer herself.” Geng (2024) also echoes this viewpoint.
@@ 4 The View from Everywhere: Frame representations are all that exist
In Bamonti and Gomes (2024b), a map m called external diffeomorphism is introduced to relate the two frames $\{\phi_{r}^{(I)}\}$ and $\{\phi_{b}^{(I)}\}$, which function analogously to a coordinate transformation. Unlike an ‘ordinary diffeomorphism’, an external diffeomorphism acts directly on the already constructed local, gauge-invariant observables, changing frames and getting us to a different and new observable.
@@
As I will further stress below, it is important to note that there can indeed be frame-independent, explicit observables, however these quantities are inherently non-local in nature (see Table 1 and my notion of Relality below).
@@
Furthermore, it is worth noting that the map m and the possibility of inter-frame communication is a frame-independent fact. This does not contradict the principles of the View From Everywhere, which do not exclude the possibility of frame-independent facts but only the possibility of frame-free facts. This last observation is fundamental to the introduction of the notion of Relality, which is based on the View From Everywhere. [^30] This flexibility of the View From Everywhere also protects it from the so-called ‘iteration of relativisation’ attacks reported in the literature, e.g. in Riedel (2024), from which similar relational approaches are said to suffer.
Relality
@@
The existence of the external diffeomorphism m enables advocates of the ”View from Everywhere” to define an equivalence class $[g_{IJ}]:=\{g_{IJ}(\phi_{r}),g_{IJ}(\phi_{b}),\dots\}$, which differs fundamentally from the equivalence class $[g_{ab}]:=\{g_{ab},(d^{*}g)_{ab},\dots\}$ used in the View from Nowhere. Unlike the latter, the equivalence class $[g_{IJ}]$ within the View from Everywhere supervenes on the ensemble of all the frame-dependent quantities. Each of these quantities form the a bona-fide physical reality, rather than existing as perspectives of a ‘more complete’, frame-free structure.
Coherently, $[g_{IJ}]$ is a frame-independent object, as it is expressed across all possible reference frames. So differently from $[g_{ab}]$, it is not a frame-free object. In formal terms: the View from Everywhere retains a one-to-one correspondence between the equivalence class $[g_{IJ}]$ and the orbit $\mathcal{O}_{g}:=\{g_{IJ}(\phi_{r}),g_{IJ}(\phi_{b}),\cdots\}$. However, unlike the View from Nowhere, $[g_{IJ}]$ in the View from Everywhere does not exist independently of the frame-specific representations within the space of models. Essentially, $[g_{IJ}]$ and $\mathcal{O}_{g}$ coincide. See Figure 3.
@@
Figure 3: The set $[M]$ of equivalence classes is part of the space of models itself. Moreover, it is the set of relational observables that can be understood as a fibre, generated by the group $Diff(\mathbb{R}^{4})$, whose group elements are external diffeomorphisms m.
@@
If for the View from Nowhere the equivalence class $[g_{ab}]$ represented a shared ontology, for the View from Everywhere, the equivalence class $[g_{IJ}]$ only represents merely the presence of a shared vocabulary. This distinction underscores the divergence between the two views: the View from Everywhere rejects the existence of a frame-free structure, focusing instead on the relational and frame-dependent nature of observables. Consequently, the term ’perspective’ becomes problematic in this context, as it implies an underlying, autonomous reality that can be described from multiple viewpoints—an assumption the View from Everywhere does not adopt.
@@
What conception of reality emerges from this framework? Within the View from Everywhere, each of the frame-dependent physical situations constitutes a fully-fledged reality. The collection of such local realities give rise to what I call Relality, which however is not an ontologically independent structure of individual frame-dependent representations: it is only a name that indicates the possibility of inter-translatability of frame-dependent and local fundamental realities.
The concept of Relality is intriguing and deserves to be analysed further. Importantly, this concept helps me to emphasise that reality is fundamentally local and relational.
Drawing on the Kantian perspective used to interpret the View from Nowhere, I argue that the View from Everywhere fundamentally rejects the notion of a shared, implicit noumenal reality underlying its phenomenological manifestations. [^31] This reference to phenomena is in no way related to French (2023)’s phenomenological approach. Here, I make no reference to the cognitive role or knowledge of conscious observers. An observer is understood as any (even non-conscious) physical system and ‘what appears to him’ is understood simply as what he enters into relation with through some physical interaction. Thus, the phenomenon is understood as the relational empirical content as encoded within the partial/complete observables framework. According to the View from Everywhere, there is no ’hidden’ structure beyond these frame-specific representations; each gauge-invariant quantity constitutes the entirety of what fundamentally exists. Accordingly, the two observables $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ do not describe two distinct phenomenal perspectives of a shared, ontologically independent noumenal reality represented by $[g_{ab}]$. Instead, each constitutes a local reality and there is nothing else. This is because the concept of Relality, formalised by $[g_{IJ}]$, does not represent a shared ontology: it is not an ontological structure ‘beyond’ and in addition to realities $\{g_{IJ}(\phi_{r}),g_{IJ}(\phi_{b}),\dots\}$. The definition of $[g_{IJ}]$ only ensures the inter-translatability of each local reality. [^32] One should not confuse this definition of Relality with the so-called first-person-plural view typical of epistemic perspectivalism. For in that view, the common vision of the community is broader and more complete than the individual perspectives on that vision. By contrast, according to my concept of Relality, the whole is not ‘something more’ than the individual parts. Relality only encompasses the interconnectedness of complete realities, but there is no collective, total vision to be discovered or achieved: neither epistemically nor ontologically. In this regard, the notion of Relality should also not be understood as confirming or disconfirming the legitimate presence of frame-independent, explicit and non-local observables. It only concerns local, relational observables. This clarification is important: it also prevents the possible attack on the frame-independent m-map according to which the exchange of information via connections between frames cannot be defined with respect to a frame-independent collective structure ($[g_{IJ}]$ precisely), which presupposes the existence of such connections in order to be defined (see e.g. Adlam, 2024b, §4.3). The map m is frame-independent in the sense that it is defined relative to any observer: there is no need for a collective structure relative to which it is defined. All of this supports the abandonment of the concept of ‘perspective’.
@@
I argue that the interpretive framework of the View from Everywhere challenges the core assumptions of the Sophistication approach. Specifically, I will demonstrate that the View from Everywhere offers greater ontological parsimony—minimising unnecessary metaphysical commitments—and provides a clearer, more perspicuous characterisation of the ontology underlying relational observables. Consequently, I contend that Sophistication is neither the ultimate nor the most effective framework for interpreting relational observables.
@@
Therefore, supporting View from Nowhere does not imply the removal of redundant gauge degrees of freedom. Rather, it only involves attributing ontological importance to the gauge-invariant observables.
@@
First, the breaking of the diffeomorphic freedom of theory is precluded by the existence of the external diffeomorphism m. This map relates different reference frame choices333333m serves as the analogue of the transition map between sections of a principal bundle, given by $\mathrm{\bf{t}}_{\sigma\sigma^{\prime}}(\varphi)=f_{\sigma^{\prime}}(\varphi)^{-1}f_{\sigma}(\varphi)$ (see Gomes (2024a))., ensuring GR to remain m-covariant (see also Bamonti and Gomes (2024a)).
Second, the notion of Relality implies that all the available information resides within each frame. Consequently, there is no ‘total’ information that could be lost when a specific reference frame is selected. This may also alleviate the concerns raised in Wallace (2024) that the choice of certain reference frames (some gauge-fixings) does not preclude the physical features of the system from varying under a gauge transformation, leading to a loss of genuine information and casting doubt on whether they can really be classified as gauge-fixing at all (see ‘The Incompleteness Reason’, ivi, p.13). [^34] To be precise, Wallace studies the so-called ‘unitary gauge’ in electrodynamics. I argue that this choice of gauge corresponds to the use of uncoupled reference frames. In other words, the physical features that change are the features of the physical system considered isolated from the environment (Wallace, 2022a, b). This isolation procedure coincides with the dynamic uncoupling of the reference frame. The “loss of genuine information” referred to by Wallace, therefore, is not really genuine because it is a consequence of this approximate procedure. Within the View from Nowhere, on the other hand, there is the loss of bona-fide genuine information as a result of the use of coupled reference frames, which identify distinct gauge-invariant observables. The View from Nowhere alleviates this more serious problem.